OpenStudy (anonymous):

Find the equation of a line parallel to x-2y=6 and has the same x- intercept as 4x+3y=-12

OpenStudy (anonymous):

Well, the cool thing is that a parallel line will have the same slope.

OpenStudy (asadkarim7):

x=3 slope = -1/2

OpenStudy (aroub):

yes do it and the parallel lines have the same slope

OpenStudy (anonymous):

yeah, i know so wats the equation?

OpenStudy (anonymous):

So, can you find the slope and slope intercept of each of those equations?

OpenStudy (anonymous):

x-2y=6 in slope intercept form =? 4x+3y=-12 in slope intercept form =?